The force exerted by a drop of water of pressure P=2.3 x 105Pa on a glass wall of surface area S=0.0042m2 is?


1
Expert's answer
2021-07-27T01:47:01-0400

Gives

"P=2.3\times10^5 pa\\S=0.0042m^2"

We know that

Pressure

"P=\frac{F}{A}"

"F=P\times A"

"F=2.3\times10^5\times0.0042=966N"